Circular No. 6647

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cl IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (R = 18.3) on V and R CCD images taken on May 2 by S. Sabine on the Mount Stromlo 1.27m telescope (+ Macho Camera). SN 1997cl is located near Abell 3577 at α = 13^{h}54^{m}18^{s}.71, δ = −27^{o}44′27′′.9 (equinox 2000.0). A nearby star has position end figures 15^{s}.02, 43′56′′.3. SN 1997cl was also present on frames taken on Apr. 5 (R = 18.0), but it is not visible on images taken on 1996 Aug. 5 (limiting mag R = 20.7).

X0726–260 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, and A. Peele, GSFC and National Research Council report: “A scan over the position of the transient Be star x-ray source X0726–260 on May 5 with the PCA on RXTE detected emission at a level of ∼ 10 mCrab, showing the source to be active. An examination of the ASM light curve of X0726–260, obtained between 1996 Jan. 5 and 1997 May 1, shows a mean flux of ∼ 2 mCrab, and the Fourier Transform of this curve has its strongest peak at a period of 34.5 days and an amplitude of ∼ 1.6 mCrab. A sine-wave fit at this period yields maximum predicted flux on May 4, consistent with the scan detection. With this ephemeris the next predicted maximum is on June 7, and further observations are encouraged at that time.”

GRS 1915+105 R. Remillard and E. Morgan, Massachusetts Institute of Technology

(MIT); J. Swank, Goddard Space Flight Center (GSFC); and the RXTE ASM instrument team at MIT and GSFC report: “The RXTE All Sky Monitor has detected bright flares from GRS 1915+105, ending a long interval of steady hard-x-ray emission at 200–300 mCrab (2–12 keV) that had been observed since early Dec. 1996. Recent flares reached 0.69 Crab on 1997 May 6.75 UT and 0.81 Crab on May 8.09. The flares are evident on timescales at least as short as hours, and there is a moderately softer x-ray spectrum at 2–12 keV. We encourage radio and infrared observations.”

1997 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green